Crew 211 Sol 1 Summary Report 22-APR-2019
Summary Title: First day on Mars: first experiments and EVA
Author’s name: Carl-Henrik Dahlqvist
Mission Status: Nominal
Sol Activity Summary: We start the day with a fitness session of 15 minutes which definitely helped some crew members to wake up. After the breakfast, we spent one hour and a half to clean the Hab. Most of the crew members have then started their respective experiments (see crew scientist report) while the crew astronomer started to work on the solar telescope. After the lunch, we all worked on our experiments until 4 pm when half of the crew went in EVA. As the UCL to Mars crew likes challenges, we will try tonight to prepare the first Martian Sushis and end the day with a meeting to discuss about today’s achievements and tomorrow’s objectives.
Look Ahead Plan: An EVA is planned in the morning to explore the Candor Chasma. Half of the crew will stay at the base to work on their experiments. Depending on the weather, the crew astronomer will try to get the first pictures of the sun and observed some fainter structures as the visual observations didn’t reveal any large perturbations in the solar atmosphere. The afternoon will again be devoted to the experiments.
Anomalies in work: Water tank pump leaking, water heater and heater in the Hab lab not working
Weather: Hot, sunny and fresh breeze.
Crew Physical Status: Good except for one of the crew member who twisted his ankle.
EVA: We have had our first EVA this afternoon which lasted for two hours. We traveled to the hills on the left of the Cow Dung Road between the Robert’S Rock Garden and Zurbin’s head to take a few soil samples to be used by Chloé, the GreenHab officer, for one of her two experiments. The EVA went very well and the sunny weather allowed us to take some beautiful pictures of the Martian environment surrounding the MDRS.
Reports to be filed: Journalist Report, GreenHab Report , HSO report, Operations Report,Astronomy Report, Scientist Report, EVA report, EVA Request, Daily Photo Report
Support Requested: None